# Linear cost and exponentially convergent approximation of Gaussian Matérn processes

### v1

@@ Appendix B Collected proofs
where $\widetilde{C}$ is a constant that only depends on $\kappa,\sigma$ and $\alpha$. Now, observe that since $\alpha>1$, we have that $\lfloor\alpha\rfloor\geq 1$. Further, note that $\|(1+w^{2})^{s/2-\lfloor\alpha\rfloor}\|_{L_{2}(\mathbb{R})}<\infty$ if, and only if, $2\lfloor\alpha\rfloor-s>\nicefrac{{1}}{{2}}$. In particular, since $\lfloor\alpha\rfloor\geq 1$ we can find $s>\nicefrac{{1}}{{2}}$ such that $\|(1+w^{2})^{s/2-\lfloor\alpha\rfloor}\|_{L_{2}(\mathbb{R})}<\infty$. Therefore, if $\alpha>1$, we can find $s>\nicefrac{{1}}{{2}}$ such that $\|\varrho^{\alpha}-\varrho_{m}^{\alpha}\|_{H^{s}(\mathbb{R})}\leq\breve{C}e^{-2\pi\sqrt{\{\alpha\}m}},$ where the constant $\breve{C}$ depends on $s>\nicefrac{{1}}{{2}}$, $\kappa$, $\sigma$ and $\alpha$. Now, from Sobolev embedding, see, e.g., (McLean, 2000, Theorem 3.26), if $s>\nicefrac{{1}}{{2}}$, we have that there exists $C_{sob}>0$, depending only on $s$, such that for every $x\in\mathbb{R}$, $|\varrho^{\alpha}(x)-\varrho_{m}^{\alpha}(x)|\leq C_{sob}\|\varrho^{\alpha}-\varrho_{m}^{\alpha}\|_{H^{s}(\mathbb{R})},$ so that
$\sup_{x\in\mathbb{R}}|\varrho^{\alpha}(x)-\varrho_{m}^{\alpha}(x)|\leq C_{sob}\|\varrho^{\alpha}-\varrho_{m}^{\alpha}\|_{H^{s}(\mathbb{R})}\leq C_{sob}\breve{C}e^{-2\pi\sqrt{\{\alpha\}m}}.$
@@
Therefore, $\|r_{m}^{\alpha}-r^{\alpha}\|_{L_{\infty}(I\times I)}\leq\sup_{x\in\mathbb{R}}|\varrho^{\alpha}(x)-\varrho_{m}^{\alpha}(x)|\leq C_{sob}\breve{C}e^{-2\pi\sqrt{\{\alpha\}m}}.$ This proves (6).
Proof B.4 (of Proposition 2.1).
Note that
@@
$f_{m}^{\alpha}(w)=$ $A\sigma^{2}\kappa^{-2\alpha}\bigg{[}\frac{k\kappa^{2\lfloor\alpha\rfloor}}{(\kappa^{2}+w^{2})^{\lfloor\alpha\rfloor}}$ (16)
$~{}+\sum_{i=1}^{m}r_{i}\kappa^{2\lfloor\alpha\rfloor+2}\bigg{(}\sum_{j=1}^{\lfloor\alpha\rfloor}\frac{1}{(\kappa^{2}p)^{\lfloor\alpha\rfloor}(\kappa^{2}(1-p_{i})+w^{2})}-\frac{1}{(\kappa^{2}p)^{\lfloor\alpha\rfloor+1-j}(\kappa^{2}+w^{2})^{j}}\bigg{)}\bigg{]}.$
Finally, the explicit expression for the covariance function (4) thus directly follows by taking the inverse Fourier transform of spectral density (16) and using the linearity of the inverse Fourier transform.
Proof B.5 (of Proposition 3.1).
We begin by recalling that if $v(\cdot)$ is a stationary Gaussian process on $\mathbb{R}$ with spectral density $h$, then by (Pitt, 1971, Theorem 10.1), the process $v$ will be a Markov process of order $p$, $p\in\mathbb{N}$, on $\mathbb{R}$, if, and only if, the spectral density of $v$, namely $h$, is a reciprocal of a polynomial of degree $2p$.
The spectral density of the process $u_{0}$ is $f_{m,0}^{\alpha}$ and the spectral density of $u_{i}$, $i=1,\ldots,m$, is $f_{m,i}^{\alpha}$, where $f_{m,0}^{\alpha}$ and $f_{m,i}^{\alpha}$ were defined in (3). Moreover, observe that $f_{m,0}^{\alpha}$ is a reciprocal of a polynomial with degree $2\lfloor\alpha\rfloor$, and $f_{m,i}^{\alpha}$, $i=1,\ldots,m$, are reciprocals of polynomials of degree $2(\lfloor\alpha\rfloor+1)$, thus $u_{0}$ is a markov process of order $\lfloor\alpha\rfloor$ and $u_{j},j=1,\ldots,m$ are Markov processes of order $\lfloor\alpha\rfloor+1$. Now, it is well-known, see Whittle (1963) and (Lindgren, 2012, Theorem 4.7 and 4.8), that $u_{0}$ also solves
$\begin{cases}k(\kappa+\partial_{x})(\kappa^{2}-\Delta)^{\frac{\lfloor\alpha\rfloor-1}{2}}u_{0}=\mathcal{W}&\text{when}~{}\lfloor\alpha\rfloor~{}\text{is odd},\\ k(\kappa^{2}-\Delta)^{\frac{\lfloor\alpha\rfloor}{2}}u_{0}=\mathcal{W}&\text{when}~{}\lfloor\alpha\rfloor~{}\text{is even},\end{cases}$
@@
$\begin{cases}r_{i}(\kappa\sqrt{1-p_{i}}+\partial_{x})(\kappa+\partial_{x})(\kappa^{2}-\Delta)^{\frac{\lfloor\alpha\rfloor-1}{2}}u_{i}=\mathcal{W},&\text{when}~{}\lfloor\alpha\rfloor~{}\text{is odd},\\ r_{i}(\kappa\sqrt{1-p_{i}}+\partial_{x})(\kappa^{2}-\Delta)^{\nicefrac{{\lfloor\alpha\rfloor}}{{2}}}u_{i}=\mathcal{W},&\text{when}~{}\lfloor\alpha\rfloor~{}\text{is even},\end{cases}$
on interval $I$. Now, since each process is stationary, we start with the stationary distribution, i.e., ${\boldsymbol{\mathrm{u}}}_{0}=[u_{0}(0),u^{\prime}_{0}(0),\dotsc,u_{0}^{(\lfloor\alpha\rfloor)}(0)]^{\top}$ and ${\boldsymbol{\mathrm{u}}}_{i}=[u_{i}(0),u_{i}^{\prime}(0),\dotsc,u_{i}^{(\lceil\alpha\rceil)}(0)]^{\top}$, $i>0$, follow the stationary distribution, and thus, the restrictions of these processes from $\mathbb{R}$ to interval $I$ will also be Markov. Therefore, the tridiagonal structure of the precision matrix follows from the fact that if we consider the open set $O_{j}=(j-1,j+1)$, $j\geq 2$, $j\in\mathbb{N}$, contained in interval $I$, then, in view of the Markov property we have that for each $i=0,\dotsc,m$, ${\boldsymbol{\mathrm{u}}}_{i,j}\perp{\boldsymbol{\mathrm{u}}}_{i,k}|{\boldsymbol{\mathrm{u}}}_{i,j-1},{\boldsymbol{\mathrm{u}}}_{i,j+1}$ for $k\in I\backslash\{j-1,j,j+1\}$, where ${\boldsymbol{\mathrm{u}}}_{0,j}=(u_{0}(w_{j}),\dotsc,u_{0}^{(\lfloor\alpha\rfloor)}(w_{j}))$ and ${\boldsymbol{\mathrm{u}}}_{i,j}=(u_{i}(w_{j}),\dotsc,u_{i}^{(\lceil\alpha\rceil)}(w_{j}))$, $i=1,2,\dotsc,m$, $w_{j}\in O_{j}$. Finally, by using this conditional independences and standard techniques (see, for example, the computations of Rue & Held (2005) for conditional autoregressive models), we obtain the expressions for the local precision matrices ${\boldsymbol{\mathrm{Q}}}_{i,j}$, $i,j=1,\ldots,n$.
Proof B.6 (of Proposition 3.3.).
First, recall that $N=n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right).$ According to Algorithm 2.9 in Rue & Held (2005), the computational cost of factorizing an $n\times n$ band matrix with bandwidth $p$ is $n(p^{2}+3p)$, and solving a linear system via back-substitution requires $2np$ floating-point operations. Since ${\boldsymbol{\mathrm{Q}}}$ is an $N\times N$ matrix with bandwidth $12\lfloor\alpha\rfloor+$ 1, the total number of floating-point operations required for the Cholesky factorization is $n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right)\left((2\lfloor\alpha\rfloor+1)^{2}+3(2\lfloor\alpha\rfloor+1)\right).$ Similarly, the floating-point operations required for solving the linear system are $2n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right)(2\lfloor\alpha\rfloor+1).$
Proof B.7 (of Proposition 3.5).
By reordering $\bar{{\boldsymbol{\mathrm{U}}}}$ by location, i.e.
@@
we have that ${\boldsymbol{\mathrm{Q}}}_{\bar{{\boldsymbol{\mathrm{U}}}}|{\boldsymbol{\mathrm{y}}}}$ is an $N\times N$ matrix, with bandwidth $\lceil\alpha\rceil(m+1)$. Following the same strategy as in the proof of Proposition 3.3 gives that the Cholesky factor can be computed in

### v2

@@ Appendix B Collected proofs
Now, because $\alpha>1$, we have that $\lfloor\alpha\rfloor\geq 1$. Further, $\|(1+w^{2})^{s/2-\lfloor\alpha\rfloor}\|_{L_{2}(\mathbb{R})}<\infty$ if, and only if, $2\lfloor\alpha\rfloor-s>\nicefrac{{1}}{{2}}$. In particular, since $\lfloor\alpha\rfloor\geq 1$ we can find $s_{\alpha}>\nicefrac{{1}}{{2}}$ such that ${\|(1+w^{2})^{s_{\alpha}/2-\lfloor\alpha\rfloor}\|_{L_{2}(\mathbb{R})}<\infty}$. Therefore, if $\alpha>1$, we can find $s_{\alpha}>\nicefrac{{1}}{{2}}$ such that
$\|\varrho_{\alpha}-\varrho_{m,\alpha}\|_{H^{s_{\alpha}}(\mathbb{R})}\leq A\sigma^{2}\kappa^{-2\alpha}\breve{C}_{\alpha}C_{\{\alpha\}}e^{-2\pi\sqrt{\{\alpha\}m}},$
where the constant $\breve{C}_{\alpha}$ depends on $s_{\alpha}$, which only depends on $\alpha$. Now, from Sobolev embedding, see, e.g., (McLean, 2000, Theorem 3.26), since $s_{\alpha}>\nicefrac{{1}}{{2}}$, we have that there exists $C_{sob,\alpha}>0$, depending only on $s_{\alpha}$, thus only on $\alpha$, such that for every $x\in\mathbb{R}$, $|\varrho_{\alpha}(x)-\varrho_{m,\alpha}(x)|\leq C_{sob,\alpha}\|\varrho_{\alpha}-\varrho_{m,\alpha}\|_{H^{s_{\alpha}}(\mathbb{R})},$ so that
$\sup_{x\in\mathbb{R}}|\varrho_{\alpha}(x)-\varrho_{m,\alpha}(x)|\leq C_{sob,\alpha}\|\varrho_{\alpha}-\varrho_{m,\alpha}\|_{H^{s}(\mathbb{R})}\leq A\sigma^{2}\kappa^{-2\alpha}C_{sob,\alpha}\breve{C}_{\alpha}C_{\{\alpha\}}e^{-2\pi\sqrt{\{\alpha\}m}}.$
Therefore,
@@
$\|r_{m,\alpha}-r_{\alpha}\|_{L_{\infty}(I\times I)}\leq\sup_{x\in\mathbb{R}}|\varrho_{\alpha}(x)-\varrho_{m,\alpha}(x)|\leq A\sigma^{2}\kappa^{-2\alpha}K_{\alpha}e^{-2\pi\sqrt{\{\alpha\}m}},$
where $K_{\alpha}>0$ is a constant that only depends on $\alpha$. This proves (7).
We now move to the proof of Proposition 3. To this end, we start by recalling the following result (Saff and Stahl, 1995, Lemma 2.1):
Lemma 14
Let $10<\alpha<$ 1.
  1. 1.
    The best uniform rational approximation $R_{m}^{*}(x)$ of the function $x^{\alpha}$ on $[0,1]$, among rational functions whose numerator has degree at most $m$, is such that both the numerator and denominator have degree exactly $m$.
  2. 2.
    All $m$ zeros $\tilde{z}_{1},\dots,\tilde{z}_{m}$ and poles $\tilde{p}_{1},\dots,\tilde{p}_{m}$ of $R_{m}^{*}(\cdot)$ lie on the negative half-axis and are interlacing; i.e., with an appropriate numbering,
    $0>\tilde{z}_{1}>\tilde{p}_{1}>\tilde{z}_{2}>\tilde{p}_{2}>\dots>\tilde{z}_{m}>\tilde{p}_{m}>-\infty.$ (20)
We are now in a position to prove Proposition 3.
Proof [of Proposition 3] Define $\widetilde{P}_{m}(x)=\sum_{i=0}^{m}a_{i}x^{i}$ and $\widetilde{Q}_{m}(x)=\sum_{i=0}^{m}b_{i}x^{i}$ so that the best rational approximation of $x^{\alpha}$ on $[0,1]$ is $R_{m}^{*}(x)=\widetilde{P}_{m}(x)/\widetilde{Q}_{m}(x)$. Then, by Lemma 14, $\widetilde{P}_{m}(x)=m_{\widetilde{P}}(x-\tilde{z}_{1})(x-\tilde{z}_{2})\cdots(x-\tilde{z}_{m}),$ and $\widetilde{Q}_{m}(x)=m_{\widetilde{Q}}(x-\tilde{p}_{1})(x-\tilde{p}_{2})\cdots(x-\tilde{p}_{m}),$ where $m_{\widetilde{P}}$ and $m_{\widetilde{Q}}$ are constants and $\{\tilde{z}_{i}\}_{i=1}^{m},\{\tilde{p}_{i}\}_{i=1}^{m}$ satisfy (20). Let $\operatorname{sgn}(x)=|x|/x$ denote the sign of $x\neq 0$. Since $\tilde{z}_{i},\tilde{p}_{i}<0$ for $i=1,\ldots,m$, we have that for every $x\in(0,1)$, $\operatorname{sgn}(\widetilde{P}_{m}(x))=\operatorname{sgn}(m_{\widetilde{P}})$ and $\operatorname{sgn}(\widetilde{Q}_{m}(x))=\operatorname{sgn}(m_{\widetilde{Q}})$. Since $x^{\{\alpha\}}>0$ for $x\in(0,1)$, $\widetilde{P}_{m}(x)$ and $\widetilde{Q}_{m}(x)$ must have the same sign for $x\in(0,1)$, so that
$\operatorname{sgn}(m_{\widetilde{P}})=\operatorname{sgn}(m_{\widetilde{Q}}).$ (21)
Now, observe that $P_{m}(x)=x^{m}\widetilde{P}_{m}(x^{-1})$ and $Q_{m}(x)=x^{m}\widetilde{Q}_{m}(x^{-1}).$ Thus,
$P_{m}(x)$ $=x^{m}\widetilde{P}_{m}(x^{-1})=m_{\widetilde{P}}(1-x\tilde{z}_{1})\cdots(1-x\tilde{z}_{m})$
$=m_{\widetilde{P}}\tilde{z}_{1}\cdots\tilde{z}_{m}(-1)^{m}\left(x-\frac{1}{\tilde{z}_{1}}\right)\cdots\left(x-\frac{1}{\tilde{z}_{m}}\right)$
$=m_{P}(x-z_{1})\cdots(x-z_{m}),$ (22)
where $m_{P}=m_{\widetilde{P}}\tilde{z}_{1}\cdots\tilde{z}_{m}(-1)^{m}$ and $z_{i}=1/\tilde{z}_{i}$ for $i=1,\ldots,m$. Similarly, we have that
$Q_{m}(x)$ $=x^{m}\widetilde{Q}_{m}(x^{-1})=m_{\widetilde{Q}}(1-x\tilde{p}_{1})\cdots(1-x\tilde{p}_{m})$
$=m_{\widetilde{Q}}\tilde{p}_{1}\cdots\tilde{p}_{m}(-1)^{m}\left(x-\frac{1}{\tilde{p}_{1}}\right)\cdots\left(x-\frac{1}{\tilde{p}_{m}}\right)$
$=m_{Q}(x-p_{1})\cdots(x-p_{m}),$ (23)
where $m_{Q}=m_{\widetilde{Q}}\tilde{p}_{1}\cdots\tilde{p}_{m}(-1)^{m}$ and $p_{i}=1/\tilde{p}_{i}$ for $i=1,\ldots,m$. Now, observe that
$\operatorname{sgn}(m_{P})=(-1)^{m}\operatorname{sgn}(\tilde{z}_{1}\cdots\tilde{z}_{m})\operatorname{sgn}(m_{\widetilde{P}})=(-1)^{2m}\operatorname{sgn}(m_{\widetilde{P}})=\operatorname{sgn}(m_{\widetilde{P}}),$
where we used (20). Similarly, we have that $\operatorname{sgn}(m_{Q})=\operatorname{sgn}(m_{\widetilde{Q}})$ so that (21) implies that
$\operatorname{sgn}(m_{P})=\operatorname{sgn}(m_{Q}).$ (24)
Furthermore, (20) implies
$0>p_{n}>z_{n}>p_{n-1}>z_{n-1}>\dots>p_{1}>z_{1}>-\infty.$ (25)
Therefore, $P_{m}(\cdot)$ and $Q_{m}(\cdot)$ have no common roots and do not have multiple roots. In particular, the function $R_{m}(x)=P_{m}(x)/Q_{m}(x)$ can be decomposed in partial fractions as
$R_{m}(x)=k+\sum_{i=1}^{m}\frac{c_{i}}{x-p_{i}}.$ (26)
Observe that we already have by (25) that $p_{i}<0$ for $i=1,\ldots,m$. It remains to show that $k>0$ and $c_{i}>0$ for $i=1,\ldots,m$. Let us start by handling $k$. Note that by (24)
$k=\lim_{x\to\infty}R_{m}(x)=\lim_{x\to\infty}\frac{P_{m}(x)}{Q_{m}(x)}=\lim_{x\to\infty}\frac{m_{P}\left(1-\frac{z_{1}}{x}\right)\cdots\left(1-\frac{z_{m}}{x}\right)}{m_{Q}\left(1-\frac{p_{1}}{x}\right)\cdots\left(1-\frac{p_{m}}{x}\right)}=\frac{m_{P}}{m_{Q}}>0.$
To conclude the proof it remains to show that $c_{i}>0$ for $i=1,\ldots,m$. To this end, observe that $p_{1},\dots,p_{m}$ are simple poles of $R_{m}(x)$ and that $c_{i}$ is the residual of $R_{m}$ associated to $p_{i}$. Indeed, by (26), we have that
$\lim_{x\to p_{i}}(x-p_{i})R_{m}(x)=\lim_{x\to p_{i}}\left[c_{i}+k(x-p_{i})+(x-p_{i})\sum_{j\neq i}\frac{c_{j}}{x-p_{j}}\right]=c_{i}.$
Furthermore, recall that $Q_{m}(p_{i})=0$ for $i=1,\ldots,m$, so that
$c_{i}$ $=\lim_{x\to p_{i}}(x-p_{i})\frac{P_{m}(x)}{Q_{m}(x)}=P_{m}(p_{i})\lim_{x\to p_{i}}\frac{x-p_{i}}{Q_{m}(x)}=P_{m}(p_{i})\lim_{x\to p_{i}}\frac{x-p_{i}}{Q_{m}(x)-Q_{m}(p_{i})}$
$=P_{m}(p_{i})\lim_{x\to p_{i}}\frac{1}{\frac{Q_{m}(x)-Q_{m}(p_{i})}{x-p_{i}}}=\frac{P_{m}(p_{i})}{Q_{m}^{\prime}(p_{i})}.$
So $c_{i}=P_{m}(p_{i})/Q_{m}^{\prime}(p_{i})>0$ for $i=1,\ldots,m$. Let us first study the sign of $P_{m}(p_{i})$. By (25), we have $m-i$ values in $\{z_{i}\}_{i=1}^{m}$ that are larger than $p_{i}$, so that
$\operatorname{sgn}(P_{m}(p_{i}))=\operatorname{sgn}\Bigl{(}m_{P}\prod_{j=1}^{m}(p_{i}-z_{j})\Bigr{)}=\operatorname{sgn}(m_{P})(-1)^{m-i}.$
Let us now study the sign of $Q_{m}^{\prime}(p_{i})$. We have that
$Q_{m}^{\prime}(x)=m_{Q}\sum_{j=1}^{m}\prod_{k\neq j}(x-p_{k})\quad\text{so that}\quad Q_{m}^{\prime}(p_{i})=m_{Q}\prod_{j\neq i}(p_{i}-p_{j}).$
By (25) we have $m-i$ values in $\{p_{j}\}_{j=1}^{m}$ that are strictly larger than $p_{i}$, so that
$\operatorname{sgn}(Q_{m}^{\prime}(p_{i}))=\operatorname{sgn}\Bigl{(}m_{Q}\prod_{j\neq i}(p_{i}-p_{j})\Bigr{)}=\operatorname{sgn}(m_{Q})(-1)^{m-i}.$
Therefore, by (24), we have that
$\operatorname{sgn}(c_{i})=\frac{\operatorname{sgn}(P_{m}(p_{i}))}{\operatorname{sgn}(Q_{m}^{\prime}(p_{i}))}=\frac{\operatorname{sgn}(m_{P})(-1)^{m-i}}{\operatorname{sgn}(m_{Q})(-1)^{m-i}}=1,$
for $i=1,\ldots,m$. This shows that $c_{i}>0$ for $i=1,\ldots,m$ and concludes the proof.
Proof [of Proposition 5] Note that
@@
$f_{m,\alpha}(w)=$ $A\sigma^{2}\kappa^{-2\alpha}\bigg{[}\frac{k\kappa^{2\lfloor\alpha\rfloor}}{(\kappa^{2}+w^{2})^{\lfloor\alpha\rfloor}}$
$+\sum_{i=1}^{m}c_{i}\kappa^{2\lfloor\alpha\rfloor+2}\bigg{(}\sum_{j=1}^{\lfloor\alpha\rfloor}\frac{1}{(\kappa^{2}p)^{\lfloor\alpha\rfloor}(\kappa^{2}(1-p_{i})+w^{2})}-\frac{1}{(\kappa^{2}p)^{\lfloor\alpha\rfloor+1-j}(\kappa^{2}+w^{2})^{j}}\bigg{)}\bigg{]}.$
Finally, the explicit expression for the covariance function (9) thus directly follows by taking the inverse Fourier transform of spectral density $f_{m,\alpha}$ and using the linearity of the inverse Fourier transform.
Proof [of Proposition 6] We begin by recalling that if $v(\cdot)$ is a stationary Gaussian process on $\mathbb{R}$ with spectral density $h$, then by (Pitt, 1971, Theorem 10.1), the process $v$ will be a Markov process of order $p$, $p\in\mathbb{N}$, on $\mathbb{R}$, if, and only if, the spectral density of $v$, namely $h$, is a reciprocal of a polynomial of degree $2p$.
The spectral density of the process $u_{0}$ is $f_{m,0,\alpha}$ and the spectral density of $u_{i}$, where $i=1,\ldots,m$, is $f_{m,i,\alpha}$, where $f_{m,0,\alpha}$ and $f_{m,i,\alpha}$ were defined in (8). Moreover, observe that $f_{m,0,\alpha}$ is a reciprocal of a polynomial with degree $2\lfloor\alpha\rfloor$, and $f_{m,i,\alpha}$, $i=1,\ldots,m$, are reciprocals of polynomials of degree $2(\lfloor\alpha\rfloor+1)$, thus $u_{0}$ is a markov process of order $\lfloor\alpha\rfloor$ and $u_{j},j=1,\ldots,m$ are Markov processes of order $\lfloor\alpha\rfloor+1$. Now, it is well-known, see Whittle (1963) and (Lindgren, 2012, Theorem 4.7 and 4.8), that $u_{0}$ also solves
$\begin{cases}k(\kappa+\partial_{x})(\kappa^{2}-\Delta)^{\frac{\lfloor\alpha\rfloor-1}{2}}u_{0}=\mathcal{W}&\text{when}\leavevmode\nobreak\ \lfloor\alpha\rfloor\leavevmode\nobreak\ \text{is odd},\\ k(\kappa^{2}-\Delta)^{\frac{\lfloor\alpha\rfloor}{2}}u_{0}=\mathcal{W}&\text{when}\leavevmode\nobreak\ \lfloor\alpha\rfloor\leavevmode\nobreak\ \text{is even},\end{cases}$
@@
$\begin{cases}c_{i}(\kappa\sqrt{1-p_{i}}+\partial_{x})(\kappa+\partial_{x})(\kappa^{2}-\Delta)^{\frac{\lfloor\alpha\rfloor-1}{2}}u_{i}=\mathcal{W},&\text{when}\leavevmode\nobreak\ \lfloor\alpha\rfloor\leavevmode\nobreak\ \text{is odd},\\ c_{i}(\kappa\sqrt{1-p_{i}}+\partial_{x})(\kappa^{2}-\Delta)^{\nicefrac{{\lfloor\alpha\rfloor}}{{2}}}u_{i}=\mathcal{W},&\text{when}\leavevmode\nobreak\ \lfloor\alpha\rfloor\leavevmode\nobreak\ \text{is even},\end{cases}$
on interval $I$. Now, since each process is stationary, we start with the stationary distribution, i.e., ${\boldsymbol{\mathrm{u}}}_{0}=[u_{0}(0),u^{\prime}_{0}(0),\dotsc,u_{0}^{(\lfloor\alpha\rfloor)}(0)]^{\top}$ and ${\boldsymbol{\mathrm{u}}}_{i}=[u_{i}(0),u_{i}^{\prime}(0),\dotsc,u_{i}^{(\lceil\alpha\rceil)}(0)]^{\top}$, $i>0$, follow the stationary distribution, and thus, the restrictions of these processes from $\mathbb{R}$ to interval $I$ will also be Markov. Therefore, the tridiagonal structure of the precision matrix follows from the fact that if we consider the open set $O_{j}=(j-1,j+1)$, $j\geq 2$, $j\in\mathbb{N}$, contained in interval $I$, then, in view of the Markov property we have that for each $i=0,\dotsc,m$, ${\boldsymbol{\mathrm{u}}}_{i,j}\perp{\boldsymbol{\mathrm{u}}}_{i,k}|{\boldsymbol{\mathrm{u}}}_{i,j-1},{\boldsymbol{\mathrm{u}}}_{i,j+1}$ for $k\in I\backslash\{j-1,j,j+1\}$, where ${\boldsymbol{\mathrm{u}}}_{0,j}=(u_{0}(w_{j}),\dotsc,u_{0}^{(\lfloor\alpha\rfloor)}(w_{j}))$ and ${\boldsymbol{\mathrm{u}}}_{i,j}=(u_{i}(w_{j}),\dotsc,u_{i}^{(\lceil\alpha\rceil)}(w_{j}))$, $i=1,2,\dotsc,m$, $w_{j}\in O_{j}$. Finally, by using this conditional independences and standard techniques (see, for example, the computations of Rue and Held (2005) for conditional autoregressive models), we obtain the expressions for the local precision matrices ${\boldsymbol{\mathrm{Q}}}_{i,j}$, $i,j=1,\ldots,n$.
Proof [of Proposition 8.] First, recall that $N=n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right).$ According to Algorithm 2.9 in Rue and Held (2005), the computational cost of factorizing an $n\times n$ band matrix with bandwidth $p$ is $n(p^{2}+3p)$, and solving a linear system via back-substitution requires $2np$ floating-point operations. Since ${\boldsymbol{\mathrm{Q}}}$ is an $N\times N$ matrix with bandwidth $12\lfloor\alpha\rfloor+$ 1, the total number of floating-point operations required for the Cholesky factorization is $n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right)\left((2\lfloor\alpha\rfloor+1)^{2}+3(2\lfloor\alpha\rfloor+1)\right).$ Similarly, the floating-point operations required for solving the linear system are $2n\left(m\lceil\alpha\rceil+\max(\lfloor\alpha\rfloor,1)\right)(2\lfloor\alpha\rfloor+1).$
Proof [of Proposition 10] By reordering $\bar{{\boldsymbol{\mathrm{U}}}}$ by location, i.e.
@@
we have that ${\boldsymbol{\mathrm{Q}}}_{\bar{{\boldsymbol{\mathrm{U}}}}|{\boldsymbol{\mathrm{y}}}}$ is an $N\times N$ matrix, with bandwidth $\lceil\alpha\rceil(m+1)$. Following the same strategy as in the proof of Proposition 8 gives that the Cholesky factor can be computed in